OpenStudy (anonymous):

a center wants to set up experimental garden. it has 320 meters of fencing to enclose a rectangle area of 6300 square meters. find the length and width of the rectangle as follows. what are the demnsions of the rectangle

OpenStudy (anonymous):

2L + 2W = P = 320 -> L + W = 160 -> L = 160 - W LW = 6300 -> (160 - W)(W) = 6300 W^2 - 160W + 6300 = 0 W = 70, and L = 90

OpenStudy (anonymous):

All good now @lexigolds9 ?

OpenStudy (anonymous):

The key is getting the perimeter equation in terms of one variable and then using that one variable for the area equation and getting a quadratic equation in ""W".

OpenStudy (anonymous):

yes thank you. need help with another tho. can u help?

OpenStudy (anonymous):

LW gives the area, but that is in 2 variables. But once you make the substitution, you have an equation in one variable. Sure , go ahead.

OpenStudy (anonymous):

5x^2=3+3x x=3+squareroot of 69/10 , 3-squareroot of 69/10 now i need to find the appropirate solutions for it

OpenStudy (anonymous):

for x

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recogntion! @lexigolds9

OpenStudy (anonymous):

are you not gonna help?

OpenStudy (anonymous):

Yes, I am going to help, but I need time to think.

OpenStudy (anonymous):

ok ill keep it open

OpenStudy (anonymous):

It looks to me like you just did solve it. You are saying that you need to find solutions, but you just listed them quite accurately. What's the problem?

OpenStudy (anonymous):

If you are looking for answers that don't have the radical sign, then you can also state the solutions as: 1.130662386 and -0.530662386 but those decimals go on forever. Your representation of the solution is sufficient and 100% accurate.

OpenStudy (anonymous):

I'm not sure if this is the type of answer you are looking for, but there was nothing wrong with your answer. All I did was get rid of the radical sign. But your answer is actually the preferred way to state the answer since sqrt(69) is an irrational number. Hope that helps. If, not, I'll stick around to see what it is you want instead, if anything.

OpenStudy (anonymous):

it is. the teacher said round.

OpenStudy (anonymous):

Ok, good, because that's all I could see as for what you wanted. All that is left is to ask rounding to what decimal place?

OpenStudy (anonymous):

but those arent right. just tried it.

OpenStudy (anonymous):

Hold on, I'll check also.

OpenStudy (anonymous):

3 deci places tho

OpenStudy (anonymous):

Yes, my answers are correct. I checked them, they work. If you want them to 3 decimal places, the answers would be: 1.131 -0.531

OpenStudy (anonymous):

yay! those worked!

OpenStudy (anonymous):

When you work with either the long numbers or the rounded numbers, you will get an answer that would also need to be rounded and that's why it looks funny, but the answers are right, don't worry about that.

OpenStudy (anonymous):

Good work! You actually did a lot of the work, so you get a medal.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

You are very welcome. Sorry I did not get back to you right away, but I am a slow typer!

OpenStudy (anonymous):

anyway you could help me rounmd this one.. 5+ squareroot of 109/6 and 5- squareroot of 109/6

OpenStudy (anonymous):

Ok, this is the last one because I have to go then and help a senior citizen who is ill.

OpenStudy (anonymous):

ok! thank you.

OpenStudy (anonymous):

If you round these to 3 decimal places, you should get: 2.573 and -0.907